# MF-LAL: Drug Compound Generation Using Multi-Fidelity Latent Space Active Learning

### v1

@@ Abstract
Current generative models for drug discovery primarily use molecular docking as an oracle to guide the generation of active compounds. However, such models are often not useful in practice because even compounds with high docking scores do not consistently show experimental activity. More accurate methods for activity prediction exist, such as molecular dynamics based binding free energy calculations, but they are too computationally expensive to use in a generative model. To address this challenge, we propose Multi-Fidelity Latent space Active Learning (MF-LAL), a generative modeling framework that integrates a set of oracles with varying cost-accuracy tradeoffs. Unlike previous approaches that separately learn the surrogate model and generative model, MF-LAL combines the generative and multi-fidelity surrogate models into a single framework, allowing for more accurate activity prediction and higher quality samples. We train MF-LAL with a novel active learning algorithm to further reduce computational cost. Our experiments on two disease-relevant proteins show that MF-LAL produces compounds with significantly better binding free energy scores than other single and multi-fidelity approaches.
@@ 1 Introduction
Due to its speed, molecular docking is the prevalent physics-based method to evaluate novel compounds by generative models (Eckmann et al., 2022; Jeon & Kim, 2020; Lee et al., 2023; Noh et al., 2022; Fu et al., 2022; Peng et al., 2022; Guan et al., 2023a; b). However, docking is a relatively poor predictor of activity (Pinzi & Rastelli, 2019; Handa et al., 2023; Coley et al., 2020; Feng et al., 2022), so it would be desirable to apply more accurate binding free energy calculation techniques (Pinzi & Rastelli, 2019; Feng et al., 2022). These techniques, based on molecular dynamics simulations, are currently considered the most reliable approach to predict affinity (Moore et al., 2023; Cournia et al., 2021). However, they have not been used by generative models due to their high computational cost (Thomas et al., 2023), with a single compound-protein pair taking hours to days to simulate on a powerful computer (Wan et al., 2020). Thus, neither docking nor binding free energy techniques alone can guide the real-world application of generative models.
@@
Multi-fidelity surrogate models aim to fuse multiple data sources as oracles spanning a range of accuracy and cost (Fernández-Godino et al., 2016).They are frequently learned using an active learning approach, where a model selects or generates queries that it is most uncertain about to send to a chosen oracle (Ren et al., 2021). The results from the oracle are then added to the training data of the model. We will focus on “query synthesis” approaches (Angluin, 1988), where the model generates its own queries to send to the oracles, speeding up learning compared to approaches that query oracles with samples from a fixed candidate set.
@@
Combining docking (low fidelity) and binding free energy (high fidelity) using multi-fidelity surrogate models holds promise to make generative models more practical. Yet, the use of multi-fidelity methods in drug discovery has been limited. Prior work from Hernandez-Garcia et al. (2023) uses a generative model to generate query compounds with high acquisition function values computed by a separate multi-fidelity surrogate model. However, since we want to generate query compounds to send to oracles at multiple fidelity levels, the distribution of optimal query compounds may differ across fidelities. A separate generative model is not aware of such differences across fidelity levels, hence it cannot send queries to the multi-fidelity oracles efficiently.
@@
We aim to address the problem of multi-fidelity generation with Multi-Fidelity Latent space Active Learning (MF-LAL), an integrated framework for compound generation using multi-fidelity active learning. Instead of separating the generative model and surrogate model, we perform surrogate modeling and generation together at each fidelity level using a sequence of hierarchical latent spaces. This improves the quality of generated queries because there is a separate latent space and decoder specialized for each fidelity, and improves surrogate modeling and inter-fidelity information passing because each latent space can be organized for predicting at just that level. We use both docking and binding free energy methods as oracles in our multi-fidelity environment to achieve a favorable trade-off between cost and accuracy. In summary,
@@
  • we introduce a novel multi-fidelity surrogate and generative modeling framework, MF-LAL, which integrates data from multiple fidelity levels to generate high-quality samples at the highest fidelity (binding free energy).
@@
  • we evaluate MF-LAL and baseline methods on a real-world problem setting involving optimizing the binding free energy of compounds against two disease-relevant human proteins, and find that MF-LAL generates compounds with significantly better scores than baselines.
@@ 2.1 Molecular generative models
One approach to guide generative models in optimizing compound binding affinity is to use an oracle for compound evaluation. This oracle can be applied to reinforcement learning (Jeon & Kim, 2020; Fu et al., 2022; Mazuz et al., 2023), VAEs (Eckmann et al., 2022; Noh et al., 2022), genetic algorithms (Spiegel & Durrant, 2020; Fu et al., 2022), diffusion models (Lee et al., 2023; Hoogeboom et al., 2022; Wu et al., 2024), or other generative frameworks Zhu et al. (2024). All of them use docking software, such as AutoDock (Morris et al., 2009), as the oracle, because it is the only reasonably fast option. However, docking is known to be inaccurate (Pinzi & Rastelli, 2019), and compounds with high docking scores do not consistently show experimental activity (Handa et al., 2023; Coley et al., 2020; Feng et al., 2022).
Molecular dynamics-based binding free energy calculations are much more accurate than docking (Moore et al., 2023; Cournia et al., 2021), but have not yet been applied to de novo generative drug design due to their high computational cost (Thomas et al., 2023). While Ghanakota et al. (2020) use binding free energy calculations in combination with a molecular generative model, they focus on the optimization of an existing known lead compound. This allows them to rely on much cheaper relative binding free energy calculations, as opposed to the absolute binding free energy (ABFE) calculations needed for de novo design (Cournia et al., 2017).
Structure-based generative models are trained on 3D structures of protein-ligand pairs, and aim to predict a 3D ligand that fits in a given protein pocket with high binding affinity. Techniques include autoregressive generation (Peng et al., 2022) and diffusion modeling (Guan et al., 2023a; b). Despite not needing an oracle like docking during the generation process, the generated compounds are still evaluated with docking as a post-processing step. This means structure-based generative models do not avoid the issue of inaccurate binding affinity prediction.
@@ 2.2 Multi-fidelity surrogate modeling
Multi-fidelity modeling methods aim to fuse multiple data sources of variable accuracy and cost (Fernández-Godino et al., 2016), and are widely used in scientific fields for surrogate modeling and uncertainty quantification (Brevault et al., 2020). A popular choice of surrogate model is a Gaussian process (GP), which performs well in low data settings and produces well-calibrated uncertainty estimates (Brevault et al., 2020). One such technique to apply GPs to multi-fidelity modeling is described by Wu et al. (2020), where a downsampling kernel is used to output predictions at each fidelity level. Other multi-fidelity surrogate modeling approaches utilize neural processes (Wang & Lin, 2020; Wu et al., 2022; 2023; Niu et al., 2024) and ordinary differential equations (Li et al., 2022) as an alternative to GPs.
Multi-fidelity models are frequently trained in an active learning fashion, where one uses an estimate of a model’s uncertainty to most efficiently acquire more datapoints from an oracle (Ren et al., 2021). In the multi-fidelity setting, this means iteratively querying across both the sampling space and each different fidelity oracle (Li et al., 2020; Hernandez-Garcia et al., 2023). Traditional active learning involves selecting from a fixed candidate set with the highest acquisition function value to query oracles with, which limits the training set to only existing samples. It also limits how much the model can learn with each query, since the maximally informative sample may not be present in the candidate set. Query synthesis approaches (Angluin, 1988) have been proposed to avoid this problem by using a generative model to generate new queries. Hernandez-Garcia et al. (2023) have applied these ideas to drug discovery problems by training a generative model to optimize the acquisition function computed by a separate multi-fidelity surrogate model, which does not take into account the different distributions of optimal query compounds at each fidelity.
@@ 3 MF-LAL
Figure 1: Overview of Multi-Fidelity Latent space Active Learning (MF-LAL).
We introduce Multi-Fidelity Latent space Active Learning (MF-LAL), an integrated framework for compound generation using multi-fidelity Bayesian active learning. An overview of our framework is shown in Figure 1. We encode molecules into a hierarchy of latent spaces (left panel), one for each fidelity, and learn surrogates that predict the oracle output based on the latent vectors (middle panel). These oracles are used to reverse optimize in the latent spaces to generate new compounds with high predicted scores. The generated molecules are fed to an oracle at a chosen fidelity in an active learning loop, the output of which is used to re-train the latent representation and surrogate models (right panel). After training, we use the surrogates to reverse optimize in the highest fidelity latent space and generate compounds with high property scores at the highest fidelity. See Appendix A for details of our model and a diagram of the network architecture.
@@ 3.1 Learning multi-fidelity latent representations
A multi-fidelity environment consists of a set of oracles $\{f_{1},\dots,f_{k},\dots,f_{K}\}$ that predict a property of interest, where the accuracy and cost of the predictions increase with the fidelity level $k$. We have a multi-fidelity dataset $\mathcal{D}$ consisting of $K$ distinct sets of molecules, one for each fidelity, $\mathcal{D}=\{\{x_{1}^{(i)},y_{1}^{(i)}\}_{i=1}^{N_{1}},\dots,\{x_{K}^{(i)},y_{K}^{(i)}\}_{i=1}^{N_{K}}\}$. Each $y_{k}^{(i)}=f_{k}(x_{k}^{(i)})$ is the result from querying oracle $k$ with molecule $x_{k}^{(i)}$. The molecules are drawn from unknown distributions $p^{*}_{1},\dots,p^{*}_{K}$. We aim to approximate these distributions using generative models $p_{\theta_{1}},\dots,p_{\theta_{K}}$ with parameters $\theta_{1},\dots,\theta_{K}$. Note that $p_{1}^{*}\neq\dots\neq p_{K}^{*}$, meaning we must learn separate generative models for each fidelity level, as opposed to previous approaches that learn a single generative model for all fidelities.
@@
In order to generate molecules with high property scores, we aim to learn differentiable surrogates $\hat{f}_{1},\dots,\hat{f}_{K}$ that approximate the oracles and use them for reverse optimization in the latent spaces. Each surrogate $\hat{f}_{k}$ maps from its corresponding latent vector $\mathbf{z}_{k}$ to an estimate of $f_{k}$. We use gradient-based optimization to find a point in a given latent space that has a high property score predicted by the surrogate, which can then be decoded to a molecule (Gómez-Bombarelli et al., 2018). The $h_{\xi_{k}}$ networks, which pass information between latent spaces, allow us to re-use information learned about the molecule’s binding properties at lower fidelities to aid in prediction at the higher fidelities without having to re-learn it using large amounts of high-fidelity data. This is because training the surrogate models organize each latent space (Tevosyan et al., 2022) for property prediction at that level, and so the latent vectors contain information about the binding properties useful for predicting the oracle output that can then be passed to higher fidelities. Additionally, the use of separate latent spaces for each fidelity level, as opposed to previous approaches that use only a single latent space shared across all levels, improves surrogate modeling performance because each latent space can be organized for prediction at just that level.
We use stochastic variational Gaussian process models (SVGPs, (Hensman et al., 2015)) with parameters $\lambda_{1},\dots,\lambda_{K}$ for the surrogates. We chose SVGPs for their speed of training and ability to train with minibatches. Specifically, $\hat{f}_{k}$ is given by $\hat{f}_{k}\sim\mathcal{GP}(m_{\lambda_{k}}(x),\Sigma_{\lambda_{k}}(x,x^{\prime}))$ where $m(x)$ and $\Sigma(x,x^{\prime})$ are the mean and covariance kernels of the SVGP. To train the model, we jointly minimize the evidence lower bound (ELBO) (Kingma & Welling, 2013) of the latent encodings and marginal log likelihood (MLL) of the GP models. Given a training molecule $x$ and associated oracle output $y$ at fidelity $k$, we minimize
[…]
where the first term is the ELBO and the second is the MLL of the GP. While the loss is only evaluated at fidelity $k$, it is backpropagated through to all lower fidelities. Additionally, in our implementation, we approximate the MLL GP loss using the ELBO (Hensman et al., 2015) for improved scalability.
@@ 3.2 Bayesian active learning for sample-efficient training
5:     $x\leftarrow$ generateHighScoringCompounds($k$, $M$, 1) (Algorithm 2)
@@
8:     if $k<K$ and $\Sigma_{\lambda_{k}}(q_{\phi}(x))<\gamma_{k}$ then
@@
Similar to Kandasamy et al. (2016), we start with only querying the oracle at the lowest fidelity level $k=1$ and increase to higher fidelities when the model’s uncertainty falls below certain thresholds. We use the posterior variance of the GP surrogate $\Sigma_{\lambda_{k}}$ to measure the model’s uncertainty. Specifically, during active learning, we repeatedly generate a latent vector $\mathbf{z}_{k}$ at fidelity $k$ that decodes to a query compound. If $\Sigma_{\lambda_{k}}(\mathbf{z}_{k})<\gamma_{k}$, where $\gamma_{k}$ is the uncertainty threshold, then we permanently increment $k$ by one for all subsequent queries. Otherwise, $k$ remains the same. Once at the highest fidelity, we keep running active learning until some computational budget is reached. We use this stepwise approach to ensure all surrogates have enough training data to make accurate predictions in high property areas of the latent spaces, thus leading to high property compounds being generated.

### v3

@@ Abstract
Current generative models for drug discovery primarily use molecular docking as an oracle to guide the generation of active compounds. However, such models are often not useful in practice because even compounds with high docking scores do not consistently show real-world experimental activity. More accurate methods for activity prediction exist, such as molecular dynamics based binding free energy calculations, but they are too computationally expensive to use in a generative model. To address this challenge, we propose Multi-Fidelity Latent space Active Learning (MF-LAL), a generative modeling framework that integrates a set of oracles with varying cost-accuracy tradeoffs. Using active learning, we train a surrogate model for each oracle and use these surrogates to guide generation of compounds with high predicted activity. Unlike previous approaches that separately learn the surrogate model and generative model, MF-LAL combines the generative and multi-fidelity surrogate models into a single framework, allowing for more accurate activity prediction and higher quality samples. Our experiments on two disease-relevant proteins show that MF-LAL produces compounds with significantly better binding free energy scores than other single and multi-fidelity approaches ($\sim$ 50% improvement in mean binding free energy score). The code is available at https://github.com/Rose-STL-Lab/MF-LAL.
Machine Learning, ICML
@@ 1 Introduction
Due to its speed, molecular docking is the prevalent physics-based method to evaluate novel compounds by generative models (Eckmann et al., 2022; Jeon & Kim, 2020; Lee et al., 2023; Noh et al., 2022; Fu et al., 2022; Peng et al., 2022; Guan et al., 2023a, b). However, docking is a relatively poor predictor of activity (Pinzi & Rastelli, 2019; Handa et al., 2023; Coley et al., 2020; Feng et al., 2022), so it would be desirable to apply more accurate binding free energy calculation techniques (Pinzi & Rastelli, 2019; Feng et al., 2022). These techniques, based on molecular dynamics simulations, are currently considered the most reliable approach to predict affinity (Moore et al., 2023; Cournia et al., 2021). However, they have not been used by generative models due to their high computational cost (Thomas et al., 2023), with a single compound-protein pair taking hours to days to simulate on a powerful computer (Wan et al., 2020). Thus, neither docking nor binding free energy techniques alone can guide the real-world application of generative models.
@@
Multi-fidelity surrogate models aim to fuse multiple data sources as oracles spanning a range of accuracy and cost (Fernández-Godino, 2023).They are frequently learned using an active learning approach, where a model selects or generates queries that it is most uncertain about to send to a chosen oracle (Ren et al., 2021). The results from the oracle are then added to the training data of the model. We will focus on “query synthesis” approaches (Angluin, 1988), where the model generates its own queries to send to the oracles, speeding up learning compared to approaches that query oracles with samples from a fixed candidate set (Guo et al., 2021; Zhu & Bento, 2017; Morand et al., 2022).
@@
Combining docking (low fidelity) and binding free energy (high fidelity) using multi-fidelity surrogate models holds promise to make generative models more practical. Yet, the use of multi-fidelity methods in drug discovery has been limited. Prior work from (Hernandez-Garcia et al., 2023) uses a generative model to generate query compounds with high acquisition function values computed by a separate multi-fidelity surrogate model. However, since we want to generate query compounds to send to oracles at multiple fidelity levels, the distribution of optimal query compounds may differ across fidelities. A separate generative model is not aware of such differences across fidelity levels, hence it cannot send queries to the multi-fidelity oracles efficiently.
@@
We aim to address the problem of multi-fidelity generation with Multi-Fidelity Latent space Active Learning (MF-LAL), an integrated framework for compound generation using multi-fidelity active learning. Instead of separating the generative model and surrogate model, we perform surrogate modeling and generation together at each fidelity level using a sequence of hierarchical latent spaces. This improves the quality of generated queries because there is a separate latent space and decoder specialized for each fidelity, and improves surrogate modeling because each latent space can be organized for predicting at just that level. Information is shared between fidelity levels using networks that map from lower to higher fidelity latent spaces. We use both docking and binding free energy methods as oracles in our multi-fidelity environment to achieve a favorable trade-off between cost and accuracy. In summary,
@@
  • We introduce a novel multi-fidelity generative modeling framework, MF-LAL, which integrates data from multiple fidelity levels to generate high-quality samples at the highest fidelity (binding free energy).
@@
  • We evaluate MF-LAL and state-of-the-art baseline methods on a real-world task of optimizing the binding free energy of compounds against two disease-relevant human proteins, and find that MF-LAL generates compounds with significantly better scores than baselines ($\sim$ 50% improvement in mean binding free energy).
@@ 2.1 Molecular generative models
One approach to guide generative models in optimizing compound binding affinity is to use an oracle for compound evaluation. This oracle can be applied to reinforcement learning (Jeon & Kim, 2020; Fu et al., 2022; Mazuz et al., 2023), VAEs (Eckmann et al., 2022; Noh et al., 2022), genetic algorithms (Spiegel & Durrant, 2020; Fu et al., 2022), diffusion models (Lee et al., 2023; Hoogeboom et al., 2022; Wu et al., 2024; Zhou et al., 2024), or other generative frameworks (Zhu et al., 2024). All of them use docking software, such as AutoDock (Morris et al., 2009), as the oracle, because it is the only reasonably fast option. However, docking is known to be inaccurate (Pinzi & Rastelli, 2019), and compounds with high docking scores do not consistently show experimental activity (Handa et al., 2023; Coley et al., 2020; Feng et al., 2022).
Molecular dynamics-based binding free energy calculations are much more accurate than docking (Moore et al., 2023; Cournia et al., 2021), but have not yet been applied to de novo generative drug design due to their high computational cost (Thomas et al., 2023). While Ghanakota et al. (2020) use binding free energy calculations in combination with a molecular generative model, they focus on the optimization of a known lead compound. This allows them to rely on much cheaper relative binding free energy calculations, as opposed to the absolute binding free energy (ABFE) calculations needed for de novo design (Cournia et al., 2017).
Structure-based generative models are trained on 3D structures of protein-ligand pairs, and aim to predict a 3D ligand that fits in a given protein pocket with high binding affinity. Techniques include autoregressive generation (Peng et al., 2022) and diffusion modeling (Guan et al., 2023a, b). Despite not needing an oracle like docking during the generation process, the generated compounds are still evaluated with docking as a post-processing step. This means structure-based generative models do not avoid the issue of inaccurate binding affinity prediction.
@@ 2.2 Multi-fidelity surrogate modeling
Multi-fidelity modeling methods aim to fuse multiple data sources of variable accuracy and cost (Fernández-Godino, 2023), and are widely used in scientific fields for surrogate modeling and uncertainty quantification (Brevault et al., 2020). A popular choice of surrogate model is a Gaussian process (GP), which performs well in low data settings and produces well-calibrated uncertainty estimates (Brevault et al., 2020). One such technique to apply GPs to multi-fidelity modeling is described by Wu et al. (2020), where a downsampling kernel is used to output predictions at each fidelity level. Other multi-fidelity surrogate modeling approaches utilize neural processes (Yating & Lin, 2020; Wu et al., 2022, 2023; Niu et al., 2024) and ordinary differential equations (Li et al., 2022b) as an alternative to GPs.
Multi-fidelity models are frequently trained in an active learning fashion, where one uses an estimate of a model’s uncertainty to most efficiently acquire more datapoints from an oracle (Ren et al., 2021). In the multi-fidelity setting, this means iteratively querying across both the sampling space and each different fidelity oracle (Li et al., 2022a; Hernandez-Garcia et al., 2023). Traditional active learning involves selecting from a fixed candidate set with the highest acquisition function value to query oracles with, which limits the training set to only existing samples. It also limits how much the model can learn with each query, since the maximally informative sample may not be present in the candidate set. Query synthesis approaches (Angluin, 1988) have been proposed to avoid this problem by using a generative model to generate new queries. Hernandez-Garcia et al. (2023) have applied these ideas to drug discovery problems by training a generative model to optimize the acquisition function computed by a separate multi-fidelity surrogate model, which does not take into account the different distributions of optimal query compounds at each fidelity.
@@ 3 MF-LAL
Figure 1: Overview of Multi-Fidelity Latent space Active Learning (MF-LAL). Left: molecules are encoded into a hierarchy of latent spaces. Middle: surrogate models predict the oracle outputs based on the latent vectors, and reverse optimization is performed in the latent spaces to generate high scoring compounds. Right: generated molecules are sent to the oracles and the results are used to re-train the latent representation and surrogate models.
We introduce Multi-Fidelity Latent space Active Learning (MF-LAL), an integrated framework for compound generation using multi-fidelity Bayesian active learning. An overview of our framework is shown in Figure 1. We encode molecules into a hierarchy of latent spaces (left panel), one for each fidelity, and learn surrogates that predict the oracle output based on the latent vectors (middle panel). These oracles are used to reverse optimize in the latent spaces to generate new compounds with high predicted scores. The generated molecules are fed to an oracle at a chosen fidelity in an active learning loop, the output of which is used to re-train the latent representation and surrogate models (right panel). After training, we use the surrogates to reverse optimize in the highest fidelity latent space and generate compounds with high property scores at the highest fidelity. See Figure 2 for a diagram of the network architecture, and Appendix A for further details of the model.
@@ 3.1 Learning multi-fidelity latent representations
A multi-fidelity environment consists of a set of oracles $\{f_{1},\dots,f_{k},\dots,f_{K}\}$ that predict a property of interest, where the accuracy and cost of the predictions increase with the fidelity level $k$. We have a multi-fidelity dataset $\mathcal{D}$ consisting of $K$ distinct sets of molecule-activity pairs, one for each fidelity, $\mathcal{D}=\{\{x_{1}^{(i)},y_{1}^{(i)}\}_{i=1}^{N_{1}},\dots,\{x_{K}^{(i)},y_{K}^{(i)}\}_{i=1}^{N_{K}}\}$. Each $y_{k}^{(i)}=f_{k}(x_{k}^{(i)})$ is the result from querying oracle $f_{k}$ with molecule $x_{k}^{(i)}$. The molecules are drawn from unknown distributions $p^{*}_{1},\dots,p^{*}_{K}$. We aim to approximate these distributions using generative models $p_{\theta_{1}},\dots,p_{\theta_{K}}$ with parameters $\theta_{1},\dots,\theta_{K}$. Note that $p_{1}^{*}\neq\dots\neq p_{K}^{*}$, meaning we must learn separate generative models for each fidelity level, as opposed to previous approaches that learn a single generative model for all fidelities.
@@
In order to generate molecules with high property scores, we aim to learn differentiable surrogates $\hat{f}_{1},\dots,\hat{f}_{K}$ that approximate the oracles and use them for reverse optimization in the latent spaces. Each surrogate $\hat{f}_{k}$ maps from its corresponding latent vector $\mathbf{z}_{k}$ to an estimate of $f_{k}$. We use gradient-based optimization to find a point in a given latent space that has a high property score predicted by the surrogate (“reverse optimization”), which can then be decoded to a molecule (Gómez-Bombarelli et al., 2018). The $h_{\xi_{k}}$ networks, which pass information between latent spaces, allow us to re-use information learned about the molecule’s binding properties at lower fidelities to aid in prediction at the higher fidelities without having to re-learn it using large amounts of high-fidelity data. This is because training the surrogate models organize each latent space (Tevosyan et al., 2022) for property prediction at that level, and so the latent vectors contain information about the binding properties useful for predicting the oracle output that can then be passed to higher fidelities. Additionally, the use of separate latent spaces for each fidelity level, as opposed to previous approaches that use only a single latent space shared across all levels, improves surrogate modeling performance because each latent space can be organized for prediction at just that level.
We use stochastic variational Gaussian process models (SVGPs, (Hensman et al., 2015)) with hyperparameters $\lambda_{1},\dots,\lambda_{K}$ (which define the mean and covariance kernels of the GP) for the surrogates. We chose SVGPs for their speed of training and ability to train with minibatches. Specifically, $\hat{f}_{k}$ is given by $\hat{f}_{k}\sim\mathcal{GP}(m_{\lambda_{k}}(x),\Sigma_{\lambda_{k}}(x,x^{\prime}))$ where $m(x)$ and $\Sigma(x,x^{\prime})$ are the mean and covariance kernels of the SVGP. To train the model, we jointly minimize the evidence lower bound (ELBO) (Kingma & Welling, 2014) of the latent encodings and marginal log likelihood (MLL) of the GP models (see Appendix A.1 for the full loss equation). While the loss is only evaluated at fidelity $k$, it is backpropagated through to all lower fidelities. Additionally, in our implementation, we approximate the MLL GP loss using the ELBO (Hensman et al., 2015) for improved scalability.
@@ 3.2 Bayesian active learning for sample-efficient training
Figure 2: Diagram of MF-LAL architecture. An input molecule is encoded into the first latent space $\mathbf{z}_{1}$ using a neural network. The networks $h_{\xi_{1}},\dots,h_{\xi_{K-1}}$ transform points in $\mathbf{z}_{1}$ to the higher fidelity latent spaces. Each latent space has an associated decoder, which reconstructs the original molecule, and a GP surrogate model for that fidelity level.
@@
5:     $x\leftarrow$ generateHighScoringMols($k$, $M$, 1) (Algorithm 2)
@@
8:     if $k<K$ and $\Sigma_{\lambda_{k}}(\mathbf{z}_{k})<\gamma_{k}$ then
@@
Similar to (Kandasamy et al., 2016), we start with only querying the oracle at the lowest fidelity level $k=1$ and increase to higher fidelities when the model’s uncertainty falls below certain thresholds. We use the posterior variance, $\Sigma_{\lambda_{k}}$, of the GP surrogate $\hat{f_{k}}$ to measure the model’s uncertainty. Specifically, during active learning, we repeatedly generate a latent vector $\mathbf{z}_{k}$ at fidelity $k$ that decodes to a query compound. If $\Sigma_{\lambda_{k}}(\mathbf{z}_{k})<\gamma_{k}$, where $\gamma_{k}$ is the uncertainty threshold (which we treat as a hyperparameter), then we permanently increment $k$ by one for all subsequent queries. Otherwise, $k$ remains the same. Once at the highest fidelity, we keep running active learning until some computational budget is reached. We use this stepwise approach to ensure all surrogates have enough training data to make accurate predictions in high property areas of the latent spaces, thus leading to high property compounds being generated.